The length of a rectangle is 1 ft. more than twice the width. The area is 55 sq. ft. Find the length and width?

Answer by checkley77(12569)

L=2W+1
W(2W+1)=55
2W^2+W-55=0
(2W+11)(W-5)=0
W-5=0
W=5 IS THE WIDTH MEASUREMENT.
L=2*5+1
L=10=1
L=11 IS THE LENGTH.
PROOF
5*11=55
55=55